OpenStudy (anonymous):

Explain the grouping method of factoring. Describe a scenario when the grouping method would be preferred over other methods and provide an example of this type of problem.

OpenStudy (anonymous):

I just need a better understanding!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Microrobot Can u help please :)

OpenStudy (anonymous):

When you group it is more organized, allowing there to be a lower chance of you making errors, it allows shows the steps done perfectly.

OpenStudy (anonymous):

Thank youuu

OpenStudy (anonymous):

Wait it says to provide an example

OpenStudy (anonymous):

So it's both needed

OpenStudy (anonymous):

yeah im gonna find an example

OpenStudy (anonymous):

like 10x^2 + 8x+ 4 ?

hero (hero):

@crystelle, Did you make that one up?

OpenStudy (anonymous):

yes!

hero (hero):

You shouldn't have....because it doesn't factor....at least not the way you need it to. If you want it to factor properly you have to make sure the value of the discriminant is square. Otherwise it won't factor.

OpenStudy (anonymous):

Oh how do I make number that will factor?

hero (hero):

Are you familiar with the discriminant?

OpenStudy (anonymous):

a little bitt..my teacher was explaing a lil

hero (hero):

\[b^2 - 4ac\]

hero (hero):

All you have to do is make sure that whatever values of a, b, c you pick if \(b^2 - 4ac\) produces a perfect square, then the quadratic will be factorable.

OpenStudy (anonymous):

Ohhh okay soo hm lemme see

hero (hero):

For example, suppose a =3 b = 2, c = -1 Then 2^2 - 4(3)(-1) = 4 + 12 = 16 And 16 is a perfect square.

hero (hero):

Therefore, you can write the following quadratic 3x^2 + 2x - 1 And it is factorable I created this one myself

hero (hero):

To factor it by grouping you would have to think of two numbers that multiplied to get -3, yet added to get 2. For this one, those two numbers would obviously be 3 and - 1 So you replace 2 with 3 - 1 3x^2 + 2x - 1 becomes 3x^2 + (3 - 1)x - 1 Then you distribute the x to get 3x^2 + 3x - 1x - 1 Then factor the first two terms to get 3x(x + 1) and the last two terms to get -1(x + 1) And as you can see x + 1 is common to both factorizations 3x(x + 1) - 1(x + 1) So factor out x + 1 : (x + 1)(3x - 1)

hero (hero):

Any questions?

OpenStudy (anonymous):

I undertand it! Thanks so muchhH!! ur really good in math

hero (hero):

I learned some of these techniques on this site.

OpenStudy (anonymous):

wow! thats pretty tightt

OpenStudy (anonymous):

hmm im a little confused on one part

OpenStudy (anonymous):

So you replace 2 with 3 - 1 3x^2 + 2x - 1 becomes 3x^2 + (3 - 1)x - 1

OpenStudy (anonymous):

@Hero

hero (hero):

Correct

OpenStudy (anonymous):

no that is what u said but i dont get the x -1 at the end

hero (hero):

2x = (3 - 1)x

hero (hero):

2 = 3 - 1

hero (hero):

All I did was replace 2 with 3 - 1 The x and -1 were already there

hero (hero):

\[3x^2 + \color\red{2x} - 1\] \[3x^2 + \color\red{(3 - 1)x} - 1\]

hero (hero):

Or if you can't see it that way then \(3x^2 + \color\green{2}x - 1\) \(3x^2 + \color\green{(3 - 1)}x - 1\)